a man is 3 times as old as his son. In 12 years time he will be twice as old as his son. How old is both of them?
Answer by macston(5194)

F=Father's age=3S; S=Son's age
F=3S
F+12=2(S+12) Substitute for F
3S+12=2(S+12)
3S+12=2S+24 Subtract 2S from each side.
S+12=24 Subtract 12 from each side.
S=12 ANSWER 1:The son is 12 years old.
F=3S=3(12)=36 years old ANSWER 2: The father is 36 years old.
CHECK:
In 12 years:
F+12 years=2(S+12 years)
36+12=2(12+12)
48 years =48 years